# Skullptor: High Fidelity 3D Head Reconstruction in Seconds with Multi-View Normal Prediction

### v1

@@ 4.3 Geometry Evaluation
We now compare our full pipeline, which produces meshes from sets of multi-view image, against other reconstruction baselines (Meshroom [23], 2DGS [29], SuGaR [24]). We first align each predicted mesh to the ground truth mesh using Procrustes analysis as described in Sec. 3.2.
Figure 4: Qualitative comparison of mesh reconstruction methods. Top: NPHM; Bottom: Multiface
For evaluation, we rasterize depth and normal maps from 12 novel camera views (unseen during reconstruction) at 512x512 resolution for each predicted and ground truth mesh. We also render the textured ground truth mesh in order to estimate a facial mask using Facer [70], to evaluate the geometry of the face region.
Given that these baseline methods are computationally intensive, we limit this comparison to a randomly chosen subset of 50 expressions/frames from each dataset (NPHM and Multiface), for a total of 100 facial reconstructions. Note that the Multiface ground truth (40-cam) and Meshroom baseline (26-cam) are both photogrammetry-based, which may lead to correlated biases for Meshroom on Multiface. For instance, note the ‘broken nose’ artifact in the Multiface photogrammetry (GT+Meshroom), which penalizes our correct reconstruction. Table 2 presents a quantitative comparison on Multiface and NPHM datasets. Our approach substantially outperforms 2DGS and SuGaR across all metrics on both datasets while being an order of magnitude faster and with signifcantly less camera views. Compared to Meshroom, our method achieves on-par geometric quality on NPHM while using less than half the views (10 vs. 23) and being an order of magnitude faster. Qualitative results (Figure 4) demonstrate that our method achieves comparable visual fidelity to Meshroom and captures fine-scale geometric details, such as wrinkles and folds, that are missed or smoothed over by 2DGS and SuGaR. Additional comparison are provided in supp. material.
@@ 4.4 Ablation Studies
Effect of Input View Count: We evaluate our pipeline’s robustness to sparse input views by comparing against Meshroom, the best performing baseline from Table 2, on NPHM using 3, 6, 10, 16, and 23 views. As shown in Figure 5 and Figure 6, our method achieves high-fidelity reconstruction even in extremely sparse configurations of only 3 cameras, while Meshroom degrades rapidly below 16 views and fails almost completely at 3 views. This demonstrates that learned geometric priors effectively compensate for reduced view coverage, enabling practical sparse-view capture setups where traditional photogrammetry fails.
@@
Impact of the Normal Estimator: To validate the importance of our multi-view normal estimator (Sec. 3.1), we replace it with monocular baselines (Sapiens 0.3B/2B, DAViD) within our full pipeline. For these monocular methods, we independently predict normals for each of the 10 input views and then run the same inverse rendering optimization to recover the geometry. Since monocular predictors process each view independently without enforcing multi-view consistency, they produce normals that can be geometrically inconsistent across viewpoints, leading to degraded reconstruction quality. As shown in Table 3, our multi-view estimator consistently produces superior geometry. Notably, while Sapiens 2B performed well on the normal comparison evaluation (Table 1), its final geometry is poor (6.76 mm error). We attribute this to its tendency to produce overly smooth normals, an observation supported by the high average normal gradient error it achieved in 1, that leads to loss of fine surface details during optimization. This discrepancy demonstrates that multi-view processing during the normal prediction stage translates to more accurate and coherent final geometry.
Table 3: Ablation study on normal estimator impact on final mesh quality (10 views).
Dataset Normal Estimator Depth (mm) $\downarrow$ Angular Err. $\downarrow$ Avg. Normal Grad. Err. $\downarrow$
@@
5 Limitations and Future Work
While our method achieves high-fidelity production-quality 3D reconstructions from sparse input views, it is designed for capture setups with controlled lighting and synchronized cameras. Strong view-dependent reflections, noisy images, and facial props can cause inaccurate normal predictions that propagate to the final geometry. Future work includes extending our framework to jointly predict normals and albedo for complete appearance capture, incorporating material and lighting estimation to enable relighting.
6 Conclusion
We presented a method for high-fidelity production-quality 3D head reconstruction from sparse multi-view images that combines the complementary strengths of learning-based and optimization-based approaches. By introducing a multi-view normal prediction model that enforces geometric consistency across viewpoints and integrating it with a custom inverse rendering optimization, our approach achieves photogrammetry-level quality from fewer than 10 camera views while being an order of magnitude faster. Our results demonstrate that learned geometric priors can effectively compensate for sparse view coverage, enabling practical high-fidelity capture setups. We believe this work represents an important step toward making professional-quality 3D facial capture more accessible in production pipelines.

### v3

@@ 4.2 Normal Evaluation
Figure 4: Qualitative comparison of mesh reconstruction methods on the NPHM (top) and Multiface (bottom) datasets.
@@ 4.3 Geometry Evaluation
We now compare our full pipeline, which produces meshes from sets of multi-view image, against other reconstruction baselines (Meshroom [23], 2DGS [29], SuGaR [24]). We first align each predicted mesh to the ground truth mesh using Procrustes analysis as described in Sec. 3.2. For evaluation, we rasterize depth and normal maps from 12 novel camera views (unseen during reconstruction) at $\times512$ resolution for each predicted and ground truth mesh. We also render the textured ground truth mesh in order to estimate a facial mask using Facer [69], to evaluate the geometry of the face region.
To ensure a representative yet computationally feasible evaluation, we limit this comparison to a randomly chosen subset of 50 expressions/frames from each dataset (NPHM and Multiface), for a total of 100 facial reconstructions. Quantitative results are summarized in Table 2.
Our approach substantially outperforms 2DGS and SuGaR across all metrics on both datasets while being an order of magnitude faster and using significantly fewer camera views. Compared to Meshroom, our method achieves on-par geometric quality on NPHM while using less than half the input views (10 vs. 23) and being an order of magnitude faster.
[…]
Qualitative results (Figure 4) demonstrate that our method achieves comparable visual fidelity to Meshroom and captures fine-scale geometric details, such as wrinkles and folds, that are missed or smoothed over by 2DGS and SuGaR. Additional comparison are provided in the supplemental material.
@@ 4.4 Ablation Studies
Effect of input view count. We evaluate our pipeline’s robustness to sparse input views by comparing against Meshroom, the best performing baseline from Table 2, on NPHM using 3, 6, 10, 16, and 23 views. As shown in Figure 5 and Figure 6, our method achieves high-fidelity reconstruction even in extremely sparse configurations of only 3 cameras, while Meshroom degrades rapidly below 16 views and fails almost completely at 3 views. This demonstrates that learned geometric priors effectively compensate for reduced view coverage, enabling practical sparse-view capture setups where traditional photogrammetry fails.
@@
Multi-view and cross-attention importance. To isolate the architectural contributions of our model, we conducted a series of ablation experiments across four distinct configurations, summarized in Tab. 3. We evaluate each setup on both normal estimation and the full Skullptor mesh reconstruction pipeline: (1) the original DAViD baseline; (2) the DAViD model fine-tuned on our dataset using only single-view inputs (DAViD_ft_mono); (3) the DAViD model trained on batches of multi-view inputs but without our cross-attention mechanism (DAViD_ft_multi); and (4) our full proposed model incorporating view-aware cross-attention with multi-view training data. The results of this analysis show that, first, simply fine-tuning alone on our dataset with monocular input (DAViD_ft_mono) does not improve performance, and can even degrade it. This confirms that the improvements observed in our full model are not merely a byproduct of domain adaptation. Secondly, training on multi-view inputs of the same subjects helps the model converge to more accurate results, demonstrating that consistent predictions are fundamental to resolving geometric ambiguities. Finally, our full model with cross-attention achieves the highest performance across all benchmarks, highlighting the structural enhancements of Skullptor’s normal prediction model.
Normal Metrics Mesh Metrics
Method Ang. $\downarrow$ Grad. $\downarrow$ Depth $\downarrow$ Ang. $\downarrow$ Grad. $\downarrow$
NPHM DAViD 7.80 0.188 2.68 5.68 0.115
DAViD_ft_mono 8.43 0.174 3.53 6.04 0.116
DAViD_ft_multi 7.13 0.176 2.87 5.98 0.113
Skullptor 7.05 0.167 2.44 5.89 0.110
Multiface DAViD 8.61 0.238 3.95 6.83 0.163
DAViD_ft_mono 8.88 0.228 4.32 7.06 0.163
DAViD_ft_multi 8.59 0.227 3.70 6.74 0.161
Skullptor 8.54 0.226 3.29 6.17 0.156
Table 3: Comparison of normal (left) and mesh (right) error.
5 Limitations and Future Work
While our method achieves high-fidelity production-quality 3D reconstructions from sparse input views, it is designed for capture setups with controlled lighting and synchronized cameras. Strong view-dependent reflections, noisy images, and facial props can cause inaccurate normal predictions that propagate to the final geometry. Future work includes extending our framework to jointly predict normals and albedo for complete appearance capture and incorporating material and lighting estimation to enable relighting.
6 Conclusion
We presented Skullptor, a method for high-fidelity production-quality 3D head reconstruction from sparse multi-view images that combines the complementary strengths of learning-based and optimization-based approaches. By introducing a multi-view normal prediction model that enforces geometric consistency across viewpoints and integrating it with a custom inverse rendering optimization, our approach achieves photogrammetry-level quality from fewer than 10 camera views while being an order of magnitude faster. Our results demonstrate that learned geometric priors can effectively compensate for sparse view coverage, enabling practical high-fidelity capture setups. We believe this work represents an important step toward making professional-quality 3D facial capture more accessible in production pipelines.
Supplementary Material
S1 Additional Results
Additional qualitative results for our method, Skullptor, including 4D sequences and extended comparisons with state-of-the-art methods for normal estimation and mesh reconstruction, are available in the supplemental webpage.
S1.1 Impact of the Normal Estimator
@@ 6 Conclusion
To validate the importance of our multi-view normal estimator (Sec. 3.1), we replace it with several monocular baselines (Sapiens 0.3B/2B, DAViD) within our full pipeline. For these monocular methods, we independently predict normals for each of the 10 input views and then run the same inverse rendering optimization to recover the geometry. Since monocular predictors process each view independently without enforcing multi-view consistency, they produce normals that can be geometrically inconsistent across viewpoints, leading to degraded reconstruction quality. As shown in Table S1, our multi-view estimator consistently produces superior geometry. Notably, while Sapiens 2B performed well on the normal comparison evaluation (Table 1), its final geometry is poor (6.76 mm error on NPHM). We attribute this to its tendency to produce overly smooth normals, an observation supported by the high average normal gradient error it achieved in Table 1, which leads to a loss of fine surface details during optimization. This discrepancy demonstrates that multi-view processing during the normal prediction stage translates to more accurate and coherent final geometry.
Dataset Normal Estimator Depth ($m$) $\downarrow$ Angular Err. $\downarrow$ Avg. Normal Grad. Err. $\downarrow$
@@
Table S1: Ablation study on normal estimator impact on final mesh quality (10 views).
To complement these quantitative findings, we provide visual comparisons in Figure S1 illustrating how different normal estimators affect the remeshing stage. Monocular predictors exhibit noticeable identity drift and loss of structural fidelity in the reconstructed mesh due to the lack of shared geometric information during view processing. In contrast, our multi-view approach maintains high-frequency details and identity consistency by leveraging cross-view features during the initial estimation phase.
Figure S1: Qualitative impact of normal estimation quality on mesh reconstruction.
S1.2 Generalization to In-the-Wild Capture
Although our primary evaluation focuses on lightstage capture setups, DAViD’s backbone provides robust priors from 300K diverse images that our fine-tuning specializes rather than overwrites. As a result, Skullptor generalizes over four distinct capture setups (TripleGangers, Multiface, NPHM, and ours) with varying camera configurations, lighting, and subjects. Additionally, we include an in-the-wild result in Fig. S2, suggesting potential for generalizing beyond controlled environments.
Figure S2: Skullptor results on 12 phone-captured images.
S2 Implementation Details For Multi-View Normal Prediction
In this section, we provide implementation details for the normal prediction (Sec. 3.1 in the main paper).
S2.1 Training Dataset
Our training dataset is curated from high-quality Triplegangers [1] facial scan assets to ensure diversity across ethnicity, age, and gender. Each scan in the original dataset was captured using a professional light stage setup comprising 55 synchronized high-end cameras, followed by a photogrammetry reconstruction pipeline with manual cleaning to ensure geometrical accuracy. Each subject is captured while performing 20 static facial expressions, resulting in 1,000 unique expression instances across the dataset. The distribution of subjects in our training dataset is given in Table S2.
Table S2: Subject distribution in the Triplegangers training dataset.
Male Female
Number of subjects 18 32
Age range 18–64 18–75
Mean age 37.2 42.8
Age << 25 9 (50%) 14 (44%)
Age 25–65 8 (44%) 12 (38%)
Age >> 65 1 (6%) 6 (19%)
Training-validation split. We reserve 5 subjects (100 expressions) for validation and use the remaining 45 subjects (900 expressions) for training. This subject-level split ensures that the model’s generalization is evaluated on entirely unseen identities.
Mesh normalization. To standardize the input geometry across subjects and expressions, we apply Procrustes analysis using 3D facial landmarks to align each scan to a normalized template facial mesh. Specifically, we center each mesh at the origin and scale it to fit within a unit sphere. This normalization ensures consistent scale and orientation across all training samples. Our foreground segmentation allows us to get the full upper part and neck, not only the face.
Rendering setup. For every subject’s expression, we render both RGB images and rasterized surface normal maps from 48 virtual cameras at $512\times 512$ resolution using nvdiffrast [37]. The rasterized normal maps are transformed into each camera’s local coordinates by multiplying them with the rotation of its model-view matrix. Camera viewpoints are randomly sampled for every expression-subject pair to maximize viewpoint diversity during training. In total, the training set consists of $43200$ images.
S2.2 Camera View Generation
As mentioned above, we generate 48 diverse camera views per expression for every subject. The camera viewpoints are selected by randomly sampling from a parameterized camera distribution that ensures comprehensive coverage of the facial region. For the 48 cameras associated with an expression, we construct a camera transformation matrix using the following random sampling of parameters as listed in Table S3. The camera transformation matrix is constructed by combining rotations around the X-axis (pitch) and Y-axis (yaw), followed by translation and scaling operations. Random sampling allows the model to learn from a wide range of viewing angles and distances.
Table S3: Camera Sampling and Projection Parameters
Camera Sampling
Pitch angle $\theta_{\text{pitch}}$ $\mathcal{U}(-35^{\circ},35^{\circ})$
Yaw angle $\theta_{\text{yaw}}$ $\mathcal{U}(-90^{\circ},90^{\circ})$
Initial radius $\mathbf{r}$ 2.0
Extra 3D translation $\mathbf{t}$ $\mathcal{U}(-0.2,0.2)^{3}$
Scale factor $z$ $\mathcal{U}(1.0,1.8)$
Projection Matrix
Image resolution $w=h$ 512 pixels
Focal lengths $f_{x}=f_{y}$ 512 pixels
Principal point $c_{x}=c_{y}$ 256 pixels
Depth range (near $n$, far $f$) 0.0010.00110001000
S2.3 Training Procedure
Each epoch iterates over all training expressions, selecting six random viewpoints from the 48 pre-rendered cameras for each expression. This sampling strategy encourages the model to integrate information across diverse view combinations, reducing the risk of overfitting to particular camera configurations.
We use the AdamW optimizer with differential learning rates: $1\times 10^{-4}$ for the newly added multi-view attention components and $1\times 10^{-5}$ for the DAViD backbone [54]. A weight decay of $1\times 10^{-4}$ is used while training the model for 100 epochs with a batch size of 6 views per sample. The model is trained on an NVIDIA Quadro RTX 8000 GPU using PyTorch, with a total training time of 82 hours on 900 training expressions (45 subjects) and 100 validation expressions (5 subjects).
S2.4 Inference
Our architecture extends the monocular foundation model [54] with a multi-view cross-attention mechanism to aggregate information across different input viewpoints in each iteration. Although the model is trained exclusively with 6-view batches, it generalizes seamlessly to arbitrary numbers of views during inference. This generalization capability stems from the inherent properties of cross-attention. Each target view attends to all available context views through learned query-key-value projections. The key and value matrices encode information from each context view into a fixed embedding dimension, which the attention mechanism aggregates regardless of the number of input views. During training with randomly sampled 6-view subsets, the model learns to flexibly combine information from diverse viewpoint configurations, making it naturally robust to different numbers of views at test time.
At inference, the model accepts any number of input views (from 3 to 26 in our experiments) without architectural modifications or retraining. The cross-attention operation dynamically adjusts to the available views, computing attention weights that reflect the geometric consistency of each viewpoint for reconstructing the target normal map.
S3 Implementation Details For Mesh Optimization
In this section, we provide implementation details for the Mesh optimization (Sec. 3.2).
For optimization, we first pre-compute the per-pixel weight matrix from the normal maps (obtained from our normal prediction model) using Eq. 9 of the main paper. The target normal maps are then masked using the foreground segmentation model from [54] to retain the complete head region, including the face, neck, and upper torso.
We use the optimization framework from Continuous Remeshing [48] to refine the mesh geometry. The optimization proceeds for 300 steps with a learning rate of 0.3. At each iteration, we rasterize the normal maps from the current mesh state using nvdiffrast (Eq. 6), compute the normal loss (Eq. 8) along with the Laplacian regularization term (Eq. 7) with $\lambda_{\mathrm{lap}}=0.1$), and update the mesh vertices. At the end of each optimization step, we apply adaptive remeshing operations from [48]. These operations consist of edge splits, collapses, and flips that dynamically adjust mesh resolution to local geometric complexity while preventing degeneracies such as collapsed faces and self-intersections.
S3.1 Initialization
For the optimization to converge successfully, the initial sphere must be positioned close to the target mesh location in 3D space. To estimate this location, we leverage the known camera extrinsics. We cast rays from each camera center through its optical axis and compute the point in 3D space that minimizes the sum of distances to all rays. This point serves as the center of the initial sphere. While this approach is most accurate when all cameras are focused on the same region, it remains robust even when camera viewing directions vary, provided the cameras roughly face a common area of interest.
During optimization, we observed that initializing from a neutral facial template rather than a sphere does not improve convergence or final quality (Figure S3). Both converge to comparable results in 300 iterations.
Figure S3: Effect of initialization on mesh optimization. Vertex color indicates local edge length (Red short large edges, blue for small edges). Top: Sphere initialization. Bottom: Template mesh initialization. Facial characteristics from the template mesh are discarded, and both initializations converge to comparable results in 300 iterations.
S3.2 Mesh Resolution
The optimization controls granularity by specifying the maximum vertex count and minimum edge length. We observed no quality gains beyond 500k vertices, suggesting resolution ceases to be a limiting factor at that scale. Skullptor’s evaluation used this vertex count, which is the same order of magnitude as 2DGS/SuGaR ($\sim$450k) and Meshroom ($\sim$250k).
S4 Evaluation Protocol on NPHM and Multiface
Because the NPHM dataset [21] provides only raw textured meshes without corresponding camera images, we generate a synthetic multi-view dataset by rendering each mesh from 23 virtual cameras. The cameras are uniformly sampled on a sphere around the head to provide full $360$ degrees of face coverage, with the head centered at the origin. We render $2048\times 2048$ color images and corresponding camera-space normal maps using NVDiffRast, without shading (texture only). These renders form the input to all reconstruction baselines. High image resolution is necessary for fair comparison, as Meshroom [23], 2DGS [29], and SuGaR [24] exhibit degraded performance at lower input resolutions.
As Multiface [28] provides original lightstage-captured images at $1334\times 2048$ resolution, we directly use the real multi-view RGB frames as input during evaluation. The dataset contains synchronized multi-view video sequences of subjects performing a variety of expressions. From the camera array, we select 26 viewpoints that provide full facial coverage and feed their captured images to all reconstruction methods.
For our method, the input images are processed to match the $512\times 512$ training resolution of our multi-view normal predictor. For NPHM, we downsample the $2048\times 2048$ rendered images. For Multiface, we resize and pad the $1334\times 2048$ captured images to obtain $512\times 512$ inputs. All baselines operate directly on the original image resolution ($2048\times 2048$ for NPHM renders and the native resolution for Multiface captures of $1334\times 2048$).
To evaluate reconstruction quality, we measure performance on novel viewpoints not used during reconstruction. We render 12 additional frontal-facing cameras, explicitly excluding any back-of-head viewpoints, to test generalization to unseen poses. For each method, we render normal and depth maps from its reconstructed mesh at $512\times 512$ from these novel viewpoints and compare them against ground-truth renderings from the corresponding ground-truth mesh under identical camera poses.
To ensure consistency across all methods, every reconstructed mesh is first aligned to a canonical coordinate system using Procrustes analysis (as described in Sec. 3.2). Depth and normal comparisons are performed after this canonical alignment for all methods, which is necessary because some baselines may output meshes in arbitrary coordinate frames.
Evaluation focuses on the frontal facial region, which is the primary area of interest for head-level fidelity. We extract per-view facial masks using Facer [69], removing hair, neck, and clothing. These masks are applied both to predictions and ground truth to ensure that metrics reflect facial geometry alone.